OpenStudy (anonymous):

Please Help just 1 Question Please Marneshia walked three over eight of a mile in three over five of an hour. What equation can be used to calculate her unit rate in miles per hour? three eighths over three fifths = five eighths mile per hour three eighths over three fifths = one and three fifths miles per hour three fifths over three eights = one and three fifths miles per hour three fifths over three eights = five eighths mile per hour

OpenStudy (anonymous):

It's the the first one (three eighths over three fifths = five eighths mile per hour).

OpenStudy (mrnood):

"per" means divided by so miles PER hour means miles/hours so miles = 3/8 hours = 3/5 so mph = 3/8 / 3/5